# Varying primordial state fractions in exo- and endothermic SIDM simulations of Milky Way-mass haloes

### v1

@@ 3.2 Excited state fraction
where delimited-⟨$\langle$..$\rangle$ … ⟩ denotes the thermal average, $\rho_{\rm gr}$ ($\rho_{\rm ex}$) denotes the density of ground (excited) state DM particles, and $\sigma_{\rm up}$ is the upscattering cross section. For simplicity, we assume that the velocities of DM particles locally obey the Maxwell-Boltzmann distribution and we obtain
@@
[…]
@@ 3.3 Halo shapes
Figure 8: We fit an ellipse to each halo at several radii and take the ratio of the middle to major axis. A ratio of 1 is spherical and a ratio of 0 is flat. The grey points show the CDM run. Each column of points is the axis ratio at that radius for each simulation halo. We see that low $\chi_{\rm init}^{2}$ haloes tend to be spherical due to early up-scattering that compresses the halo. High $\chi_{\rm init}^{2}$ haloes maintain a more oblong shape since down-scattering keeps the halo enlarged.
Figure 8 demonstrates this more quantitatively. For each halo, we fit an ellipse at different radii from 10 to 200 kpc and divide the middle axis by the major axis to measure how spherical it is. We follow the method used in Chua et al. (2021), which iteratively fits an ellipsoidal shell to the halo while keeping the semi-major axis constant. First, we fix the semi-major axis, $R$. At each step, the bin is determined by a $3\times 3$ matrix $A_{ij}$: a particle with coordinates $\vec{x}$ falls inside the bin if $0.85R\leq\sqrt{A_{ij}x_{i}x_{j}}\leq 1.15R$.
@@
Next, we diagonalize $S$ to obtain eigenvectors $v_{1},v_{2},v_{3}$ with eigenvalues $\lambda_{1},\lambda_{2},\lambda_{3}$ respectively. These determine the intermediate axis ratios $q\equiv\sqrt{\lambda_{2}/\lambda_{1}}$ and $s\equiv\sqrt{\lambda_{3}/\lambda_{1}}$. Finally, we generate a new bin with the matrix $A_{ij}^{\prime}$, which has eigenvectors $v_{1},v_{2},v_{3}$ with respective eigenvalues $1,q,s$. The iteration terminates when $q$ and $s$ change by less than $0.1\%$ between steps. For more details, see Zemp et al. (2011). A ratio of 1 is perfectly spherical, and smaller ratios are increasingly oblong. At each radius, we plot this ratio for each $\chi_{\rm init}^{2}$ in different colours.
Across all radii, haloes with a lower $\chi_{\rm init}^{2}$ are more spherical than haloes with a higher $\chi_{\rm init}^{2}$. This is consistent with previous results in O’Neil et al. (2023), where the endothermic model resulted in a more spherical halo than the exothermic model. Here, higher $\chi_{\rm init}^{2}$ haloes are dominated by exothermic reactions since the down-scattering reaction is the only available reaction when all particles are in the excited state.
Haloes do not accrete symmetrically, so it is expected that there is a certain amount of asymmetry as seen in the high $\chi_{\rm init}^{2}$ simulations. This asymmetry is wiped away by the endothermic reactions. Endothermic reactions cause particles to fall deeper into the potential well by reducing their kinetic energy during the reaction. This compresses the inner regions of the halo and causes it to become more spherical, especially when a large amount of particles are up-scattered. Even when these particles are later down-scattered and kicked further out in the halo, they cannot regain the asymmetrical orbits they previously may have had and maintain the spherical shape of the halo.
@@ 3.5 Effects on substructures
Figure 10 shows the cumulative number of subhaloes with mass $M_{\rm sub}>10^{7}{\,\rm M_{\odot}}$ within a radius $r$ to the main halo centre. In the bottom panel, we show the same quantity divided by that of the CDM simulation. In models with increasing $\chi_{\rm init}^{2}$, there are significantly less substructures. The down-scatterings increase the kinematic energy of the subhalo and make it more difficult for subhalo with small gravitational wells to stay bound. On the other hand, up-scattering decreases the velocity but the velocity dispersion in satellite haloes is generally not high enough to significantly alter the substructure in up-scattering dominant models. The simulations with higher $\chi^{2}_{\rm init}$ therefore have a much more altered substructure compared to CDM compared to simulations with a higher $\chi^{1}_{\rm init}$. However, there is still a decrease by a factor of 2 in even the $\chi_{\rm init}^{2}=0$ case since the main halo increases in central density, which increases the tidal forces and strips the subhaloes.
Figure 11 shows the cumulative subhalo mass function for the main halo in each solution (top) and the same quantity divided by that of CDM (bottom). We find that simulations with higher $\chi^{2}_{\textrm{init}}$ tend to have fewer subhaloes at all mass scales.
We can do a simple estimate of the typical subhalo mass where the heat generated from down-scattering exceeds the gravitational binding energy. For a halo with an NFW profile, the gravitational binding energy is (Mo et al., 1998)
@@
This represents the total possible heating energy to the subhalo through exothermic self-interactions. Therefore, the smallest substructure we expect to survive this heat injection is
@@
This is generally above the subhalo masses in simulations and, therefore, the numbers of subhaloes at all mass scales display significant suppression.
@@ 4 Conclusions
  • For each simulation, we fit an ellipse at various radii and compare the major and minor axes to estimate the shapes of the haloes. Figure 8 shows the minor-to-major axis ratio at several radii for each halo, so a ratio of 1 is circular and 0 is flat. Higher $\chi_{\rm init}^{2}$ consistently produces more oblong haloes across all radii with an axis ratio as low as 0.75, and low $\chi_{\rm init}^{2}$ haloes are close to spherical. This is due to early up-scattering compressing the halo such that it loses any oblong shape, and it then remains spherical throughout its evolution. With early down-scattering, the halo remains puffy and there is not a strong enough influence from scattering to compress the halo into a sphere. Figure 9 shows the anisotropy parameter $\beta$ as a function of radius for each simulation, with higher $\beta$ corresponding to more radial orbits. In the presence of self-interactions, $\beta$ decreases across all radii.
@@
The initial conditions of the halo result in significantly altered halo properties. In particular, high $\chi_{\rm init}^{2}$ results in a halo with properties consistent with those from exothermic SIDM, and low $\chi_{\rm init}^{2}$ results in properties consistent with endothermic SIDM. To accurately predict the expected halo, it is therefore important to have a solid foundation in the initial conditions of the model. In addition, several other factors would likely impact the properties of the halo that we did not consider here. The overall environment and mergers may cause certain reactions to dominate or provide an influx of new particles and substructures. The importance of this work lies in highlighting the sensitivity of the halo to early universe conditions while controlling for these other variables.
@@ Acknowledgements
Simulations and analysis were done with MIT’s Engaging cluster.

### v2

@@ 3.2 Excited state fraction
where delimited-⟨$\langle$..$\rangle$ … ⟩ denotes the thermal average, $\rho_{\rm gr}$ ($\rho_{\rm ex}$) denotes the density of ground (excited) state DM particles, and $\sigma_{\rm up}$ is the upscattering transfer cross section. For simplicity, we assume that the velocities of DM particles locally obey the Maxwell-Boltzmann distribution and we obtain
@@
Figure 6: M200,mean and Vmax for the main halo in each simulation. M200,mean decreases with $\chi^{2}_{\rm init}$ due to the kinetic energy boost from down-scattering, which can push particles out of the halo. Vmax also decreases with $\chi^{2}_{\rm init}$, reflecting trends in both halo mass and concentration. Haloes with low $\chi^{2}_{\rm init}$ experience more up-scattering, driving particles inwards.
@@ 3.3 Halo shapes
Figure 8: We fit an ellipse to each halo at several radii and find the ratio of the middle axis ($q$, top panel) and minor axis ($s$, bottom panel) to the major axis. A ratio of $s=1$ is spherical, and a ratio of $s=0$ is flat. The grey points show the CDM run. Each column of points is the axis ratio at that radius for each simulation halo. We see that low $\chi_{\rm init}^{2}$ haloes tend to be spherical due to early up-scattering that compresses the halo. High $\chi_{\rm init}^{2}$ haloes maintain a more oblong shape since down-scattering keeps the halo enlarged.
Figure 8 demonstrates this more quantitatively. For each halo, we fit an ellipse at different radii from 0.03 $R_{\rm 200,mean}$ to $R_{\rm 200,mean}$ and divide the middle and minor axes by the major axis to measure how spherical it is. We follow the method used in Chua et al. (2021), which iteratively fits an ellipsoidal shell to the halo while keeping the semi-major axis constant. First, we fix the semi-major axis, $R$. At each step, the bin is set to be within 0.1 dex of the current ellipse.
@@
Next, we obtain the eigenvectors $v_{1},v_{2},v_{3}$ with eigenvalues $\lambda_{1},\lambda_{2},\lambda_{3}$ respectively ($\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$). These determine the intermediate axis ratios $q\equiv\sqrt{\lambda_{2}/\lambda_{1}}$ and $s\equiv\sqrt{\lambda_{3}/\lambda_{1}}$. We also rotate the particles to a coordinate system with axes aligning with the eigenvectors. Finally, we generate a new bin with the new eigenvalues determining $q$ and $s$. The iteration terminates when $q$ and $s$ change by less than $0.1\%$ between steps. For more details, see Zemp et al. (2011). A ratio of $s=1$ is perfectly spherical, and smaller ratios are increasingly oblong. At each radius, we plot $q$ and $s$ for each $\chi_{\rm init}^{2}$ in different colours.
Across all radii, haloes with a lower $\chi_{\rm init}^{2}$ are more spherical than haloes with a higher $\chi_{\rm init}^{2}$. This is consistent with previous results in O’Neil et al. (2023), where the endothermic model resulted in a more spherical halo than the exothermic model. Here, higher $\chi_{\rm init}^{2}$ haloes are dominated by exothermic reactions since the down-scattering reaction is the only available reaction when all particles are in the excited state. In elastic scattering, transfer cross sections of approximately $\sigma_{T}/m_{\chi}=0.1\>{\rm cm}^{2}{\rm g}^{-1}$ drive spherical centres in the cores of galaxy clusters (e.g. Miralda-Escudé, 2002; Peter et al., 2013; Brinckmann et al., 2018; Shen et al., 2022). Here, we show that significant endothermic and elastic scattering produces similar results, but the exothermic scattering mitigates this, allowing larger elastic transfer cross sections while maintaining a certain amount of asymmetry. To fully understand the constraints in galaxy clusters, a dedicated set of simulations is needed to accurately predict how this model would behave in such environments.
Haloes do not accrete symmetrically, so it is expected that there is a certain amount of asymmetry as seen in the high $\chi_{\rm init}^{2}$ simulations. This asymmetry is wiped away by the endothermic reactions. Endothermic reactions cause particles to fall deeper into the potential well by reducing their kinetic energy during the reaction. This compresses the inner regions of the halo and causes it to become more spherical, especially when a large amount of particles are up-scattered. This is consistent with findings that low-concentration halos are generally less symmetrical than high-concentration halos (Hashimoto et al., 2007) Even when these particles are later down-scattered and kicked further out in the halo, they cannot regain the asymmetrical orbits they previously may have had and maintain the spherical shape of the halo.
@@ 3.5 Effects on substructures
Figure 10 shows the cumulative number of subhaloes with mass $M_{\rm sub}>10^{7}{\,\rm M_{\odot}}$ within a radius $r$ to the main halo centre. In the bottom panel, we show the same quantity divided by that of the CDM simulation. In models with increasing $\chi_{\rm init}^{2}$, there are significantly less substructures. The down-scatterings increase the kinematic energy of the subhalo and make it more difficult for subhalo with small gravitational wells to stay bound. On the other hand, up-scattering decreases the velocity but the velocity dispersion in satellite haloes is generally not high enough to significantly alter the substructure in up-scattering dominant models. The simulations with higher $\chi^{2}_{\rm init}$ therefore have a much more altered substructure compared to CDM compared to simulations with a higher $\chi^{1}_{\rm init}$.
However, there is still a decrease by a factor of 2 in even the $\chi_{\rm init}^{2}=0$ case since the main halo increases in central density, which increases the tidal forces and strips the subhaloes. The tidal radius $r_{t}$ of the subhalo decreases as the enclosed mass of the main halo increases
$r_{t}=r\left(\frac{m_{\rm sub}}{2M_{\rm main}\left(<r\right)}\right)^{\frac{1}{3}}$ (10)
where $m_{\rm sub}$ is the mass of the subhalo, $M_{\rm main}$ is the mass of the main halo and $r$ is the distance of the subhalo from the centre of the main halo (King, 1962; Pace et al., 2022). Thus, an increased central density of 10 will result in approximately a factor of 2 decrease in a satellite’s tidal radius.
Another important effect is the SIDM ram pressure stripping due to self-interactions between satellite and host halo DM particles (e.g. Kummer et al., 2018; Nadler et al., 2020; Jiang et al., 2021), also known as SIDM evaporation (Zeng et al., 2022). Thermalized cores generated by self-interactions can make satellites more prone to this. The relative importance of ram pressure stripping can vary when the cross-section has velocity dependence (e.g. Banerjee et al., 2020; Nadler et al., 2020; Zeng et al., 2022). In our model, the cross-section of the dominant down-scattering process decreases by orders of magnitude with increasing velocities until reaching $\sim 1000\,{\rm km}\,{\rm s}^{-1}$, which is larger by a factor of few than the maximum circular velocities of the simulated Milky Way-mass halos. The heat generated from scatterings of satellite particles dominates the heat injected from scattering host halo particles.
Figure 11 shows the cumulative subhalo mass function for the main halo in each solution (top) and the same quantity divided by that of CDM (bottom). We find that simulations with higher $\chi^{2}_{\textrm{init}}$ tend to have fewer subhaloes at all mass scales. We can do a simple estimate of the typical subhalo mass where the heat generated from down-scattering exceeds the gravitational binding energy. For a halo with an NFW profile, the gravitational binding energy is (Mo et al., 1998)
@@
This represents the maximum possible heating energy to the subhalo through exothermic self-interactions. Therefore, the smallest substructure we expect to survive this heat injection is
@@
This is generally above the subhalo masses in simulations and, therefore, the numbers of subhaloes at all mass scales display significant suppression. We note that this estimation assumes all of the energy in excited states is transformed into the kinetic energy of DM particles and should be interpreted as a limiting case. It also assumes that the down-scattering of DM particles within the satellite is the major channel of heat generation, which may not hold for other types of SIDM models or more massive host halos.
@@ 4 Conclusions
  • For each simulation, we fit an ellipse at various radii and compare the major, middle, and minor axes to estimate the shapes of the haloes. Figure 8 shows the middle-to-major and minor-to-major axis ratios ($q$ and $s$ respectively)at several radii for each halo, so a ratio of $s=1$ is circular and $q=0$ is flat. Higher $\chi_{\rm init}^{2}$ consistently produces more oblong haloes across all radii with axis ratios as low as $q=0.6$ and $s=0.55$, whereas low $\chi_{\rm init}^{2}$ haloes are close to spherical. This is due to early up-scattering compressing the halo such that it loses any oblong shape, and it then remains spherical throughout its evolution. With early down-scattering, the halo remains puffy and there is not a strong enough influence from scattering to compress the halo into a sphere. Figure 9 shows the anisotropy parameter $\beta$ as a function of radius for each simulation, with higher $\beta$ corresponding to more radial orbits. In the presence of self-interactions, $\beta$ decreases across all radii.
@@
There are several other factors that influence halo properties that we did not consider here. The overall environment and mergers may cause certain reactions to dominate or provide an influx of new particles and substructures. A dedicated set of simulations using the same early universe conditions but simulating a variety of halos would shed light on the diversity problem in particular. In addition, baryons could alter several of the properties, for example by deepening the central gravitational potential. This could cause a difference in, for example, whether subhaloes go through core collapse or not. Dissipative dark matter has been shown to accelerate gravothermal collapse (Essig et al., 2019), which mimics the behaviour of the endothermic reactions. However, the two-state dark matter particle includes exothermic reactions, which causes haloes to become more diffuse. Thus, some haloes may experience accelerated core collapse while others may not, depending on the histories of the subhaloes. Currently, most constraints exist for elastic SIDM (Adhikari et al., 2022), while other types of SIDM are not well constrained. Because the two-state model introduces several types of scattering, certain reactions will dominate in different situations. Studying how the halo properties like shape and density profile change depending on mass scale and merger history could help constrain models like this one. Larger haloes with large velocity dispersions may be more dominated by endothermic reactions since particles can more easily cross the velocity threshold. For this, it would be interesting to run a set of cluster simulations using this model.
The initial conditions of the halo result in significantly altered halo properties. In particular, high $\chi_{\rm init}^{2}$ results in a halo with properties consistent with those from exothermic SIDM, and low $\chi_{\rm init}^{2}$ results in properties consistent with endothermic SIDM. To accurately predict the expected halo, it is therefore important to have a solid foundation in the initial conditions of the model. The sensitivity of the resulting halo to the initial conditions indicate that this model can produce a range of shapes and density profiles from compact to diffuse, potentially alleviating the diversity of shapes problem. The importance of this work lies in highlighting the sensitivity of the halo to early universe conditions while controlling for these other variables.
@@ Acknowledgements
MV acknowledges support through NASA ATP 19-ATP19-0019, 19-ATP19-0020, 19-ATP19-0167, and NSF grants AST-1814053, AST-1814259, AST-1909831, AST-2007355 and AST-2107724. Simulations and analysis were done with MIT’s Engaging cluster.